# DID Connect: Authentication in TLS with Decentralized Identifiers and Verifiable Credentials

### v1

@@ VII Conclusion
This article presented DID Connect, a novel authentication scheme for TLS 1.3 that comes in two flavors, namely a basic version for the pseudo-anonymous authentication with DIDs and an advanced version for the identification with VCs. DID Connect extends TLS 1.3 with the option to let entities authenticate alternatively with artifacts of a decentrally managed digital identity, all without the need to fundamentally redesign the well approved authentication procedure of TLS. It introduces a new identification sub layer and custom extensions to the initial TLS handshake messages in the frame of permissible modifications. The identification sub layer delineates the identification-specific communication from the application data, while the custom extensions give both parties early on the opportunity to agree on common DID- and VC-specific parameters for the authentication. The adherence to the TLS 1.3 standard ensures that DID Connect can be seamlessly integrated into existing legacy systems while benefiting from a certificate-based authentication procedure that is widely used, well established and in particular thoughtfully investigated to ensure security across diverse platforms and applications. DID Connect facilitates the decentralization of the Web by enabling entities to autonomously and in a self-sovereign manner manage their identifiers and identity claims and use them for the authentication at the transport layer with TLS, without having to sacrifice the fundamental security features TLS has to offer, such as confidentiality, integrity, and authenticity.
Nevertheless, entrusting a commonly operated and governed distributed ledger with one of the essential functions of a centralized CA comes with costs. The out-of-band verification of the identifier and key binding via the ledger leads to a significant prolongation of the authentication process. But once the identification is done, the entities can fall back on the pseudo-anonymous authentication with cached DID docs, which performs even slightly better than the authentication with CA-issued X.509 certificates. Moreover, a TLS channel established with DID Connect outperforms a DIDComm link in terms of speed by a factor of more than 2 at latest after a handful of messages has been transmitted. As a next step, the VP used to carry identity claims can be equipped additionally with VCs for the authorization such like an access token, making it possible to unify and harmonize access control procedures within a single step on the transport layer. However, the exchange of VCs for access control purposes with DID Connect requires a further security analysis, as does the sharing of verification material via a common and cross-domain system in form of a distributed ledger.

### v2

@@ VII Conclusion
This article presented DID Link, a novel authentication scheme for TLS 1.3 that comes in two flavors, namely a basic version for the pseudo-anonymous authentication with DIDs and an advanced version for the identification with VCs. DID Link extends TLS 1.3 with the option to let entities authenticate alternatively with artifacts of a decentrally managed digital identity, all without the need to fundamentally redesign the well approved authentication procedure of TLS. It introduces a new identification sub layer and custom extensions to the initial TLS handshake messages in the frame of permissible modifications. The identification sub layer delineates the identification-specific communication from the application data, while the custom extensions give both parties early on the opportunity to agree on common DID- and VC-specific parameters for the authentication. The adherence to the TLS 1.3 standard ensures that DID Link can be seamlessly integrated into existing legacy systems while benefiting from a certificate-based authentication procedure that is widely used, well established and in particular thoughtfully investigated to ensure security across diverse platforms and applications. DID Link facilitates the decentralization of the Web by enabling entities to autonomously and in a self-sovereign manner manage their identifiers and identity claims and use them for the authentication at the transport layer with TLS, without having to sacrifice the fundamental security features TLS has to offer, such as confidentiality, integrity, and authenticity.
Nevertheless, entrusting a commonly operated and governed distributed ledger with one of the essential functions of a centralized CA comes with costs. The out-of-band verification of the identifier and key binding via the ledger leads to a significant prolongation of the authentication process. But once the identification is done, the entities can fall back on the pseudo-anonymous authentication with cached DID docs, which performs even slightly better than the authentication with CA-issued X.509 certificates. Moreover, a TLS channel established with DID Link outperforms a DIDComm link in terms of speed by a factor of more than 2 at latest after a handful of messages has been transmitted. As a next step, the VP used to carry identity claims can be equipped additionally with VCs for the authorization such like an access token, making it possible to unify and harmonize access control procedures within a single step on the transport layer. However, the exchange of VCs for access control purposes with DID Link requires a further security analysis, as does the sharing of verification material via a common and cross-domain system in form of a distributed ledger.